OpenStudy (anonymous):

y=2-x, x(x+y)=5-3y^2 a solution would be helpful please, no idea how these things are solved

OpenStudy (anonymous):

You know that y=2-x, so plug in 2-x instead of y in the second equation and solve.

OpenStudy (anonymous):

Hold on, lemme try

OpenStudy (anonymous):

What's (2-x)^2? Is it the long answer or like, the one with 2 numbers?

OpenStudy (anonymous):

can't do it man, help me out here

OpenStudy (mertsj):

If y = 2-x, then x = 2-y. Replace each x with 2-y

OpenStudy (anonymous):

where do I go from 2-2y=5-3y^2?

OpenStudy (mertsj):

Put it in the standard form for a quadratic equation and then solve by factoring (if it will factor) or by the quadratic formula if it will not.

OpenStudy (anonymous):

English man, English

OpenStudy (mertsj):

Rearrange the equation so that it is equal to 0. Make the y^2 term positive.

OpenStudy (mertsj):

Also, it should be 4-2y not 2-2y

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Actually, Mertsj is right, it's easier to replace x other than y.

OpenStudy (anonymous):

Eh?

OpenStudy (anonymous):

could one of you dudes just give me the answer +solution? It's REALLY late

OpenStudy (mertsj):

\[(2-y)(2-y+y)=5-3y^2\] \[(2-y)(2)=5-3y^2\] \[4-2y=5-3y^2\] \[3y^2-2y-1=0\] \[(3y+1)(y-1)=0\]

OpenStudy (mertsj):

Can you solve that for y?

OpenStudy (anonymous):

waaaait

OpenStudy (anonymous):

how did 3y^2-2y-1=0 become (3y+1)(y-1)=0 ?

OpenStudy (mertsj):

By factoring.

OpenStudy (anonymous):

what method?

OpenStudy (anonymous):

You can also use discriminant if you want.

OpenStudy (mertsj):

Too many cooks.

OpenStudy (anonymous):

mmm, well

OpenStudy (anonymous):

thanks dudes, I'll try to figure this dang thing out

OpenStudy (anonymous):

Sorry Helper Can't give more than one medal